Please help me solve this....
In September 1998 the population of the country of West Goma in millions was modeled by f(x)=16.7e^0.0015x. At the same time the population of East Goma in millions was modeled by g(x)=14.6e^0.0135x. In both formulas x is the year, where x=0 corresponds to September 1998. Assuming these trends continue, estimate the year when the population of West Goma will equal the population of East Goma.
Possible answer is 2009 but not sure

Answer by ewatrrr(24785)


Hi
Population formulas, where x is the number of years beyond 1998
16.7e^0.0015x = 14.6e^0.0135x
16.7/14.6 = e^(.0135x-.0015x) = e^.012x
ln(16.7/14.6)/.012 = x
x = 11 years (to the nearest year)
1998 + 11 = 2009, the year the populations will be equal.